# Empirically Understanding the Value of Prediction in Allocation

### v2

@@ 2 Related Work
A related literature focuses on solving allocation problems once formulated. A common approach is to learn risk scores that rank individuals for allocation [Liu et al., 2026, Wang et al., 2024], referred to as “prediction policy problems” by Kleinberg et al. [2015]. Recognizing the causal nature of these problems, counterfactual risk prediction models estimate outcomes under no intervention [Coston et al., 2020]. Allocating to those who would benefit most requires estimating heterogeneous treatment effects, drawing on a vast literature on heterogeneous effect estimation [Wager and Athey, 2018, künzel2019].
@@ 5.1 Is collecting additional information for jobseekers with limited employment history worth it? (Q2)
(a) Comparison of prediction improvements versus capacity expansion for targeting long-term unemployment risk ($\alpha=\beta=0.15$). We simulate improving predictions for jobseekers over 35 years with missing employment history versus expanding screening capacity (assuming 44 hours per slot). (a) Welfare difference as a function of RMSE reduction in the subgroup assuming that collecting missing employment information costs 11 hour per jobseeker. The yellow line marks break-even. (b) Welfare difference when prediction improvements are concentrated on subgroup members near the decision threshold ($10\%$ bandwidth).
@@
Figure 3: Equivalent capacity expansion (in caseworker hours) that would generate the same welfare gain as each level of RMSE improvement.
@@ 5.2 What is the relative value of prediction when allocating both good and harms? (Q3)
To capture this pattern, we adopt a simple utility $u(w,a)=a\cdot u(w)$ based on predicted unemployment duration $y$ where u(w)=hu(w)=-h if $w<t_{\beta}$, and $u(w)=b$ if $w\geqslant t_{\beta}$ and $b$ and $h$ represent the net benefit and harm respectively (e.g., in income prospects or employment duration). Employment offices have used similar simple targeting logics to guide intake decisions in the past [Allhutter et al., 2020].
(a) Gain from capacity expansion and uniform prediction improvement under different harm-to-benefit ratios $\tfrac{h}{b}$. All panels use $\beta=0.25$, $\alpha=0.01$, and allocate until capacity is exhausted while excluding individuals with negative expected utility.
@@
Figure 5: Ratio of welfare gains ($\beta=0.25$, $\alpha=0.01$) from prediction improvement (x-axis: RMSE reduction at $\tfrac{h}{b}=2$) versus harm reduction (y-axis: reduction in $\tfrac{h}{b}$). Heatmap values are truncated at $0.2$ and $5.0$.
@@ 6 Case Study: Poverty Targeting in Ethiopia
Figure 6: Ratio of welfare impacts under predictive targeting to welfare impacts under random targeting, as a function of program capacity (keeping benefits fixed at \$100 USD PPP per household).
@@
(a) (a) Welfare-optimizing share of budget allocated to data collection vs. transfers, as a function of program budget; (b) optimal share of data collected as a function of program budget; (c) optimal program capacity as a function of program budget.
@@ 6.1 What is the value of improving prediction when designing a new program? (Q1)
We use rvp to simulate a planner designing a new aid program with a fixed budget. With no data yet collected for program targeting, how much of the budget should be invested in PMT data collection vs. cash transfers? We assume a fixed benefit size, so the only investment in allocation available to the planner is increasing capacity (we relax this constraint in Section 6.2). We simulate aid programs with budgets up to \$2.7 billion PPP per year. [^4] This maximum budget corresponds to 100% capacity for a $100$ per year cash transfer program covering Ethiopia’s 27.3 million households if no budget were allocated fro data collection. We then identify the welfare-optimizing choice of investments in data collection for targeting vs. distribution of cash transfers.
@@ Acknowledgements
UFA acknowledges the support by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education. We are thankful to Chris Hays and Sendhil Mullainathan for insightful comments and discussions. We gratefully acknowledge the Social Sciences Computing Facility (SSCF) at UC San Diego for providing computational resources. This research utilized the Social Sciences Research and Development Environment (SSRDE) cluster.
@@ A.2 Data
We make use of a comprehensive dataset on German jobseekers. The Sample of Integrated Labour Market Biographies (SIAB) is provided via a Scientific Use File from the Research Data Centre (FDZ) of the German Federal Employment Agency (BA) at the Institute for Employment Research (IAB) [Schmucker and vom Berge, 2023a]. The SIAB is a 2% sample of all governmental labor market records in Germany, comprising approximately 60 million entries spanning 1970 to 2021. It contains information on employment status, jobseeker demographics, employment and benefits histories, and participation in active labor market programs. Further details on its content, sampling procedure, and anonymization can be found in the accompanying data report [Schmucker and vom Berge, 2023b].
@@ B.1 Background
Proxy means tests are one of the most popular approaches for targeting cash transfer programs and other social protection programs in low-income contexts where comprehensive adminsitrative data on incomes or consumption expenditures are unavailable. They are currently used to target social protection programs in at least fifty low- and middle-income countries [Barrientos, 2018]. The accuracy of proxy means tests in practice for identifying poor households have been studied in many studies across Africa Brown et al. [2018], Schnitzer and Stoeffler [2024], McBride and Nichols [2018], Asia Narayan and Yoshida [2005], and Latin America Noriega-Campero et al. [2020]. Other work has assessed the fairness of PMTs relative to other targeting approaches Noriega-Campero et al. [2020], susceptibility to intentional data misreporting and collusion Banerjee et al. [2020], and acceptability to data subjects Premand and Schnitzer [2021], Alatas et al. [2012], but little previous work has addressed the benefit of investing in PMT coverage and accuracy in comparison relative to investment in allocating resources to the cash transfer programs and other social programs targeted with PMTs. Aiken et al. [2025] provide an initial framework for such cost-benefit trade-offs using data on targeting cash transfers in Bangladesh, we expand this framework in this paper to account for additional policy levers and data availability constraints.

### v3

@@ 2 Related Work
A related literature focuses on solving allocation problems once formulated. A common approach is to learn risk scores that rank individuals for allocation [Liu et al., 2026, Wang et al., 2024], referred to as “prediction policy problems” by Kleinberg et al. [2015]. Recognizing the causal nature of these problems, counterfactual risk prediction models estimate outcomes under no intervention [Coston et al., 2020]. Allocating to those who would benefit most requires estimating heterogeneous treatment effects, drawing on a vast literature on heterogeneous effect estimation [Wager and Athey, 2018, Künzel et al., 2019].
@@ 5.1 Is collecting additional information for jobseekers with limited employment history worth it? (Q2)
(a) Comparison of prediction improvements versus capacity expansion for targeting long-term unemployment risk ($\alpha=\beta=0.15$). The subgroup consists of jobseekers over 35 with missing or unavailable employment history ($\approx 3\%$ of cases), for whom standard administrative data is less informative. Additional data collection assumed to cost 11 caseworker hour per jobseeker in this subgroup; this total budget is converted into an equivalent number of screening slots at 44 hours per slot, defining the capacity expansion baseline. Since the RMSE improvement achieved by data collection is uncertain, both panels plot the welfare difference across a continuous range of hypothetical uniform RMSE reductions, allowing the planner to read off the break-even requirement. (a) Data collection applied to all subgroup members; the yellow line marks break-even. (b) Same comparison when data collection is restricted to subgroup members near the decision threshold (10% bandwidth), where better predictions are most likely to change allocation decisions.
@@
Figure 3: Maximum caseworker hours a planner should invest in data collection to improve predictions for jobseekers over 35 with missing employment history. For each hypothetical uniform RMSE reduction, the y-axis shows the hours of capacity expansion, at 4 hours per additional screening slot, that would lead to the same welfare gain. This defines an upper bound on the cost-worthwhile investment in data collection.
@@ 5.2 What is the relative value of prediction when allocating both good and harms? (Q3)
To capture this pattern, we adopt a simple utility $u(w,a)=a\cdot g(w)$ based on predicted unemployment duration $y$ where g(w)=hg(w)=-h if $w<t_{\beta}$, and $g(w)=b$ if $w\geqslant t_{\beta}$ and $b$ and $h$ represent the net benefit and harm respectively (e.g., in income prospects or employment duration). Employment offices have used similar simple targeting logics to guide intake decisions in the past [Allhutter et al., 2020].
(a) Welfare gains from (a) capacity expansion (as $p$ increase) and (b) uniform prediction improvement for different harm-to-benefit ratios $\frac{h}{b}$, where $h$ is the cost of misallocating to a low-risk individual and $b$ is the benefit of correctly targeting a high-risk individual ($\beta=0.25$, $\alpha=0.01$). When misallocation is harmless ($\frac{h}{b}=0$), expanding capacity increases welfare; as $\frac{h}{b}$ grows, capacity expansion becomes welfare-reducing because the harm of misallocating to additional low-risk individuals outweighs the benefit of reaching additional high-risk ones. The value of prediction moves in the opposite direction. Better predictions become more valuable as misallocation grows costlier, since sharper targeting allows the agency to target support without incurring additional harm.
@@
Figure 5: Relative value of prediction improvement versus harm ($\beta=0.25$, $\alpha=0.01$, $\frac{h}{b}=2$). Each cell shows the ratio of welfare gains from a given uniform RMSE reduction (x-axis) to the welfare gain from a given reduction in $\frac{h}{b}$ (y-axis). A ratio greater than 1 (warm colors) means prediction improvement generates larger welfare gains; a ratio below 1 (cool colors) favors harm reduction. The dashed line marks the break-even contour where both levers deliver equal welfare gains. The ratio indicates how much more a planner should be willing to pay for one lever over the other. Values are truncated at $0.2$ and $5.0$ for readability.
@@ 6 Case Study: Poverty Targeting in Ethiopia
Figure 6: Value of predictive targeting relative to random allocation for cash transfer targeting in Ethiopia, as a function of program capacity $\alpha$ (benefits fixed at \$100 USD PPP per household). The y-axis shows the welfare ratio with values above 1 indicating that predictive targeting outperforms random allocation. The ratio is highest at low capacity, where scarce slots must be allocated carefully, and declines as capacity grows and targeting precision matters less. The step function utility ($\beta=0.5$) rewards reaching any poor household equally, while CRRA utility ($\rho=3$) assigns greater value to transfers reaching the poorest households, making accurate targeting more important.
@@
(a) Welfare-optimizing budget allocation for a new hypothetical cash transfer program in Ethiopia, as a function of total program budget (CRRA utility $\rho=3$, and step-function utility $\beta=0.5$). Data collection refers to poverty scorecard surveys at a median cost of \$13 USD PPP per household; capacity investment expands the share of households receiving transfers at a fixed benefit of \$100 USD PPP. (a) Optimal share of budget allocated to data collection versus transfers. (b) Resulting share of households with scorecard data collected. (c) Resulting program capacity.
@@ 6.1 What is the value of improving prediction when designing a new program? (Q1)
We use rvp to simulate a planner designing a new aid program with a fixed budget. With no data yet collected for program targeting, how much of the budget should be invested in PMT data collection vs. cash transfers? We assume a fixed benefit size, so the only investment in allocation available to the planner is increasing capacity (we relax this constraint in Section 6.2). We simulate aid programs with budgets up to \$2.7 billion PPP per year. [^4] This maximum budget corresponds to 100% capacity for a $100$ per year cash transfer program covering Ethiopia’s 27.3 million households if no budget were allocated for data collection. We then identify the welfare-optimizing choice of investments in data collection for targeting vs. distribution of cash transfers.
@@ Acknowledgements
UFA acknowledges the support by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education. We are thankful to Chris Hays and Sendhil Mullainathan for insightful comments and discussions. We thank Nanina Föhr for helpful feedback and support with the design of the paper’s illustrations. We gratefully acknowledge the Social Sciences Computing Facility (SSCF) at UC San Diego for providing computational resources. This research utilized the Social Sciences Research and Development Environment (SSRDE) cluster.
@@ A.2 Data
We make use of a comprehensive dataset on German jobseekers. The Sample of Integrated Labour Market Biographies (SIAB) is provided via a Scientific Use File from the Research Data Centre (FDZ) of the German Federal Employment Agency (BA) at the Institute for Employment Research (IAB) [Schmucker and vom Berge, 2023b]. The SIAB is a 2% sample of all governmental labor market records in Germany, comprising approximately 60 million entries spanning 1970 to 2021. It contains information on employment status, jobseeker demographics, employment and benefits histories, and participation in active labor market programs. Further details on its content, sampling procedure, and anonymization can be found in the accompanying data report [Schmucker and vom Berge, 2023a].
@@ B.1 Background
Proxy means tests are one of the most popular approaches for targeting cash transfer programs and other social protection programs in low-income contexts where comprehensive administrative data on incomes or consumption expenditures are unavailable. They are currently used to target social protection programs in at least fifty low- and middle-income countries [Barrientos, 2018]. The accuracy of proxy means tests in practice for identifying poor households have been studied in many studies across Africa Brown et al. [2018], Schnitzer and Stoeffler [2024], McBride and Nichols [2018], Asia Narayan and Yoshida [2005], and Latin America Noriega-Campero et al. [2020]. Other work has assessed the fairness of PMTs relative to other targeting approaches Noriega-Campero et al. [2020], susceptibility to intentional data misreporting and collusion Banerjee et al. [2020], and acceptability to data subjects Premand and Schnitzer [2021], Alatas et al. [2012], but little previous work has addressed the benefit of investing in PMT coverage and accuracy in comparison relative to investment in allocating resources to the cash transfer programs and other social programs targeted with PMTs. Aiken et al. [2025] provide an initial framework for such cost-benefit trade-offs using data on targeting cash transfers in Bangladesh, we expand this framework in this paper to account for additional policy levers and data availability constraints.